# Efficient Paths for Local Counterdiabatic Driving

### v1

@@ Abstract
Local counterdiabatic driving (CD) provides a feasible approach for realizing approximate reversible/adiabatic processes like quantum state preparation using only local controls and without demanding excessively long protocol times. However, in many instances getting high accuracy of such CD protocols requires engineering very complicated new controls or pulse sequences. In this work, we describe a systematic method for altering the adiabatic path by adding extra local controls along which performance of local CD protocols is enhanced. We then show that this method provides dramatic improvement in the preparation of non-trivial GHZ ground states of several different spin systems with both short-range and long-range interactions.
@@ I Introduction
However, even in the finite-size systems accessible to near-term quantum computers and modern experimental setups, the timescales required for adiabaticity are often longer than the system remains coherent and thus forbid its application. This has motivated the development of so-called “Shortcuts to adiabaticity” Guéry-Odelin et al. (2019) in which the adiabatic path may be approximately or exactly followed at faster timescales, at the expense of demanding additional control over the system.
One such technique is known as counterdiabatic (CD) driving or equivalently transitionless driving Demirplak and Rice (2003, 2005); Berry (2009); Hartmann and Lechner (2019); Wurtz and Love (2022); Chandarana et al. (2022); Schindler and Bukov (2023), where an additional counterdiabatic term is added to the Hamiltonian of the system, which exactly suppresses any transitions between states arising from a fast change of parameters. With this modified Hamiltonian the initial quantum state, which evolves according to the Schrödinger equation, will now following the instantaneous eigenstates by construction.
@@
The main purpose of this work is to propose a systematic method for adding such extra local controls to the Hamiltonian along which local CD protocols are most efficient. Schematically the idea is sketched conceptually in Figure 1. The horizontal plane represents the space of adiabatically connected ground states. By introducing extra controls we can modify the adiabatic path connecting the initial and final states (solid lines). The vertical line schematically represents an error (e.g. deviation of the fidelity from unity) resulting from the local CD driving (dashed lines). The optimal (blue) path results in a lower error. Note that while we focus on quantum state preparation this formalism can be applied to realize fast and reversible energy transfer to facilitate various thermodynamic processes Villazon et al. (2019); Gjonbalaj et al. (2022).
Figure 1: A sketch of the method for finding more efficient paths connecting an initial and the target state. We seek to improve the local CD protocol by adding extra control terms, modifying the adiabatic paths connecting these states (see also Ref. Čepaitė et al., 2023). We find that in the situations we analyzed one can drastically improve performance of the local CD protocols by following the optimal path. This is schematically illustrated by a much smaller error for an optimal path (blue) than the “naive” path (red) of the original annealing Hamiltonian.
This paper is organized as follows: in Section II we review how quantum states may be prepared adiabatically, when this fails, and how counterdiabatic driving may be employed to alleviate this. In Section III we discuss how we may employ local (approximate) counterdiabatic driving most efficiently by augmenting the underlying Hamiltonian with extra control terms. In Section IV we show how this may be applied beyond the standard short-range models, especially those connected to relevant experiments. Finally, we summarize in Section V and discuss potentially fruitful applications of these techniques beyond the contents of this paper.
@@ II Counterdiabatic driving
Counterdiabatic driving was first discovered by Demirplak & Rice Demirplak and Rice (2003, 2005) in the context of population transfer between molecular states, and independently formulated by Berry Berry (2009) and termed “transitionless driving,” although the two are entirely equivalent. This involves defining the counterdiabatic Hamiltonian $H_{CD}$ such that
$H_{CD}=H+\dot{\lambda}A_{\lambda}$ (1)
where $A_{\lambda}$ is known as the adiabatic gauge potential (AGP). This operator is responsible for transforming the instantaneous eigenstates under a change of the control parameter $\lambda$.
Formally, the AGP satisfies the following equation
$[\partial_{\lambda}H+\frac{i}{\hbar}[A_{\lambda},H],H]=0$ (2)
@@
$S_{\lambda}(A_{\lambda})=\operatorname{Tr}\left[G_{\lambda}^{2}\right];\ G_{\lambda}=\partial_{\lambda}H+\frac{i}{\hbar}[A_{\lambda},H].$ (3)
We note in passing that one can use replace $\operatorname{Tr}\left[G_{\lambda}^{2}\right]\to\operatorname{Tr}\left[\rho\,G_{\lambda}^{2}\right]$, where $\rho$ is an arbitrary stationary density matrix with respect to the Hamiltonian $H(\lambda)$ Sels and Polkovnikov (2017). We may interpret Eq. (2), which in this language reads $[G_{\lambda},H]=0$, as a statement that a well-defined AGP $A_{\lambda}$ admits the existence of a conserved operator $G_{\lambda}$ commuting with $H$.
Instead of dealing with the exact AGP, which is very nonlocal and even ill-defined in chaotic systems Pandey et al. (2020), it is convenient to find an approximate local AGP, which minimizes the action in Eq. (3) within a restricted subset of operators $A_{\lambda}$. A very convenient option is to choose this subset from the so-called Krylov space, which is obtained by a repeated action of Liouvillian on $\partial_{\lambda}H$ Claeys et al. (2019):
@@
Here $\ell$ controls the order (and therefore the locality) of the expansion, and $\alpha_{k}$ are the variational parameters found by minimizing the action (3). Formally this ansatz is exact in the limit $\ell\to\infty$ but in practice we want to restrict it to small values of $\ell$. The variational method can be further improved by using the Lanczos algorithm in the operator basis leading to an AGP expansion in an orthonormal set of Krylov operators $O_{k}$ (see Refs. Takahashi and del Campo, 2023; Bhattacharjee, 2023; Sels and Polkovnikov, 2023 for details). We give a detailed description of our Krylov space construction of the AGP in Appendix A.
@@ III Improving CD driving with extra controls
The main goal of this section and of the whole paper is to find a systematic approach to improving performance of approximate CD driving via adding extra controls and apply them to prepare nontrivial GHZ entangled states. We focus on the infinite speed limit $|\dot{\lambda}|\to\infty$, where effectively we evolve only according to the AGP as defined in Eq. (1).
@@
with periodic boundary conditions. We will start from the situation where there is a small but finite longitudinal field $0<h_{z}\ll 1$ and finite $h_{x}<1$. The first condition breaks the $Z_{2}$ symmetry of the model such that the ferromagnetic ground state for $|h_{x}|<1$ is polarized in the positive $z$-direction and the second condition implies that during the annealing protocol the system crosses a quantum phase transition at $h_{x}=1$ Sachdev (2011). At this point the gap closes and the AGP becomes an operator with an infinite range del Campo et al. (2012); Damski (2014); Kolodrubetz et al. (2017) such that local CD protocols become very inefficient.
Figure 2: A comparison between the results when performing local CD driving on the “naive” original path in red, which passes close by the critical point at $(h_{x},h_{z})=(1,0)$, and the augmented path in blue, where the critical point is avoided. The final Hamiltonian $H_{0}$ is given by Eq. (5) with $h_{x}=0.7,\;h_{z}=0.01$. The circles and squares show the results using $\ell=1$ and $\ell=2$ in Eq. (4) respectively.
@@
$\tilde{H}(\lambda)=H(\lambda)+\beta\sin(\pi\lambda)\sum_{i}\sigma_{i}^{z},$
@@ III.1 Formulation of General Problem
where $H(\lambda$) is given by Eq. (5) and the $\beta^{(n)}(\lambda)$ are some smooth functions satisfying boundary conditions $\beta^{(n)}(\lambda=0)=\beta^{(n)}(\lambda=1)=0$ so that the initial and final eigenstates of the annealing Hamiltonian are unchanged. We refer to $H_{c}$ as an extra control Hamiltonian. In the previous example, it was just an additional global $z$ field, i.e. $H_{c}=\sum_{i}\sigma_{i}^{z}$. We stress that in the infinite speed limit we follow only the AGP, so the annealing Hamiltonian plays a purely auxiliary role like e.g. in Ref. Ljubotina et al., 2022.
We then perform local CD driving for the augmented Hamiltonian $\tilde{H}(\lambda)$, hence evolving according to $\tilde{H}_{CD}(\lambda)$. The goal is to choose both $\beta^{(n)}(\lambda)$ and $H_{c}^{(n)}$ to maximize the fidelity of the final state. Finding the optimal control functions $\beta^{(n)}(\lambda)$ is the subject of quantum optimal control Glaser et al. (2015) or similar methods Bukov et al. (2018), which is not the main focus of this work. We pick the $\beta^{(n)}(\lambda)$ by a very simple optimization, following earlier work Čepaitė et al. (2023), writing it as a single harmonic term
@@
and choosing $\beta^{(n)}$ to maximize the fidelity of the final state. This is done by choosing the cost function
@@
where |$\ket{\psi_{evolved}(\beta)}$ is obtained by evolving the initial state by local $\tilde{H}_{CD}$ corresponding to $\tilde{H}(\lambda)$. We note that the fidelities can be even further improved by adding additional couplings $\beta_{k}^{(n)}$ corresponding to $k$-th harmonic of $\pi\lambda$epaitė et al. (2023). We limit ourselves only to the first one as we want to focus on the question of finding optimal $H_{c}^{(n)}$.
@@ III.2 Ansatz for Extra Controls
To proceed, let us consider a state preparation problem similar to the previous one, but with $h_{x}=h_{z}=0$, i.e. without breaking the $Z_{2}$-symmetry. The ferromagnetic ground state of this model is now a GHZ state Greenberger et al. (2007). It is defined as an equal superposition of all spins up and all spins down. This adiabatic state preparation is encoded in the Hamiltonian
@@
with the corresponding weights $\beta^{(n)}(\lambda)$. By adding such terms we clearly expand the variational manifold allowing one to gradually increase locality of both $\tilde{H}(\lambda)$ and $A_{\lambda}$. Moreover such terms can be generated using Floquet pulse sequences containing only $H_{0}$ and $H_{1}$ as was done in Ref. Claeys et al., 2019 for realizing the AGP, and is routinely done in the NMR literature Farrar and Becker (1971). As we show below for the examples we analyze here this idea leads to the dramatic improvement of local CD protocols.
It is easy to check that for the Ising model the new terms (i.e. terms which are not originally present in $H_{0}$) which appear by computing the commutators in Eq. (10) are $YY$ and $ZXZ$, where we use a common short-hand notation: $YY=\sum_{j}\sigma_{j}^{y}\sigma_{j+1}^{y}$, $ZXZ=\sum_{j}\sigma_{j}^{z}\sigma_{j+1}^{x}\sigma_{j+2}^{z}$. We can thus select $H_{c}^{(1)}=YY$ and $H_{c}^{(2)}=ZXZ$. In this way the augmented Hamiltonian (6) will cause the state to follow a genuinely new path. The resulting improvement of the final state fidelity is plotted in Figure 3. For a more detailed discussion of the optimization procedure see Appendix B.
@@
Figure 3: Improvement in the final state fidelity when prepared with local counterdiabatic driving. In red, we perform local CD driving along the original path given by Eq. (5), whereas in blue it is performed along a path where we augment Eq. (5) by $H_{c}^{(1)}=YY$ and $H_{c}^{(2)}=ZXZ$. The two sets of red and blue lines correspond to $\ell=1$ and $\ell=2$ in Eq. (4)
IV Preparing GHZ state in a Long range model.
We now move to preparation of the GHZ state in longer range Ising Hamiltonians, which are relevant to cold atom/trapped ion systems Koffel et al. (2012); Jaschke et al. (2017); Adelhardt et al. (2020). We define the annealing Hamiltonian for a long-range Ising model:

### v2

@@ Abstract
Local counterdiabatic driving (CD) provides a feasible approach for realizing approximate reversible/adiabatic processes like quantum state preparation using only local controls and without demanding excessively long protocol times. However, in many instances getting high accuracy of such CD protocols requires engineering very complicated new controls or pulse sequences. In this work, we describe a systematic method for altering the adiabatic path by adding extra local controls along which performance of local CD protocols is enhanced, both close to and far away from the adiabatic limit. We also identify an iterative procedure to improve the performance of local counterdiabatic driving further without any knowledge of the quantum wavefunction. We then show that these methods provides dramatic improvement in the preparation of non-trivial GHZ ground states of several different spin systems with both short-range and long-range interactions.
@@ I Introduction
However, even in the finite-size systems accessible to near-term quantum computers and modern experimental setups, the timescales required for adiabaticity are often longer than the system remains coherent and thus forbid its application. This has motivated the development of so-called “Shortcuts to adiabaticity” Guéry-Odelin et al. (2019) in which the adiabatic path may be approximately or exactly followed on shorter timescales, at the expense of demanding additional control over the system.
One such technique is known as counterdiabatic (CD) driving or equivalently transitionless driving Demirplak and Rice (2003, 2005); Berry (2009); Wurtz and Love (2022); Chandarana et al. (2022); Schindler and Bukov (2023); Lawrence et al. (2024), where an additional counterdiabatic term is added to the Hamiltonian of the system, which exactly suppresses any transitions between states arising from a fast change of parameters. With this modified Hamiltonian the initial quantum state, which evolves according to the Schrödinger equation, will now follow the instantaneous eigenstates by construction.
@@
The main purpose of this work is to propose a systematic method for adding such extra local controls to the Hamiltonian so that local CD protocols are most efficient. Schematically the idea is sketched conceptually in Figure 1. The horizontal plane represents the space of adiabatically connected ground states. By introducing extra controls we can modify the adiabatic path connecting the initial and final states (solid lines). The vertical line schematically represents an error (e.g. deviation of the fidelity from unity) resulting from the local CD driving (dashed lines). The optimal (blue) path results in a lower error. Note that while we focus on quantum state preparation this formalism can be applied to realize fast and reversible energy transfer to facilitate various thermodynamic processes Villazon, Polkovnikov, and Chandran (2019); Gjonbalaj, Campbell, and Polkovnikov (2022).
Figure 1: A sketch of the method for finding more efficient paths connecting an initial and the target state. We seek to improve the local CD protocol by adding extra control terms, modifying the adiabatic paths (solid lines) connecting these states (see also Ref. Čepaitė et al., 2023). We find that in the situations we analyzed one can drastically improve performance of the local CD protocols by following the optimal path. This is schematically illustrated by a much smaller error for an optimal path (blue) than the “naive” path (red) of the original annealing Hamiltonian when performing local CD driving (dashed lines).
This paper is organized as follows: in Section II we review how quantum states may be prepared adiabatically, when this fails, and how counterdiabatic driving may be employed to alleviate this. In Section III we discuss how we may employ local (approximate) counterdiabatic driving more efficiently by augmenting the underlying Hamiltonian with extra control terms. In Section IV we show how this may be applied beyond the standard short-range models, especially those connected to relevant experiments. We also identify a novel iterative procedure for further improving the local CD driving protocol in the absence of extra controls. Finally, we summarize in Section V and discuss potentially fruitful applications of these techniques beyond the contents of this paper.
@@ II Counterdiabatic driving
Counterdiabatic driving was first discovered by Demirplak & Rice Demirplak and Rice (2003, 2005) in the context of population transfer between molecular states, and independently formulated by Berry Berry (2009) and termed “transitionless driving,” although the two are entirely equivalent. This involves defining the counterdiabatic Hamiltonian $H_{\rm CD}$ such that
$H_{\rm CD}=H+\dot{\lambda}A_{\lambda}$ (1)
where $A_{\lambda}$ is known as the adiabatic gauge potential (AGP). This operator is responsible for transforming the instantaneous eigenstates under a change of the control parameter $\lambda$, and it is the term which makes the adiabatic approximation exact. There are many possible ways to encode $\lambda$ as a function of time. Following earlier work, we encode it in the following smooth way
$\lambda(t)=\sin^{2}\left(\frac{\pi}{2}\sin^{2}\left(\frac{\pi t}{2\tau}\right)\right)$ (2)
so that $\lambda(0)=0$ and $\lambda(\tau)=1$ The “speed” of the process is encoded in $\tau$. Since $\dot{\lambda}\propto 1/\tau$, when $\tau$ is small, the counterdiabatic term $\dot{\lambda}A_{\lambda}$ dominates the dynamics defined by $H_{\rm CD}$ in Eq. (1). As $\tau$ is increased, this term vanishes and the process becomes adiabatic.
Formally, the AGP satisfies the following equation (in units where $\hbar=1$)
$[\partial_{\lambda}H+i[A_{\lambda},H],H]=0$ (3)
@@
$S_{\lambda}(A_{\lambda})=\operatorname{Tr}\left[G_{\lambda}^{2}\right];\ G_{\lambda}=\partial_{\lambda}H+i[A_{\lambda},H].$ (4)
We note in passing that one can use replace $\operatorname{Tr}\left[G_{\lambda}^{2}\right]\to\operatorname{Tr}\left[\rho\,G_{\lambda}^{2}\right]$, where $\rho$ is an arbitrary stationary density matrix with respect to the Hamiltonian $H(\lambda)$ Sels and Polkovnikov (2017). By using $\rho={1\over Z}\exp{-\beta H}$ ), we can perform a finite-temperature variational optimization, which preferentially targets lower-energy eigenstates and can give better performance in some instances. In the limit $\beta\to\infty$ the average with respect to $\rho$ reduces to the ground state expectation value. In Sec. IV.2 we discuss in detail how one can further improve local CD protocols using the ground state optimization of the AGP without requiring any prior knowledge of the ground state. We may interpret Eq. (3), which in this language reads $[G_{\lambda},H]=0$, as a statement that a well-defined AGP $A_{\lambda}$ admits the existence of a conserved operator $G_{\lambda}$ commuting with $H$.
Instead of dealing with the exact AGP, which is very nonlocal and even ill-defined in chaotic systems Pandey et al. (2020), it is convenient to find an approximate local AGP, which minimizes the action in Eq. (4) within a restricted subset of operators $A_{\lambda}$. A very convenient option is to choose this subset from the so-called Krylov space, which is obtained by a repeated action of Liouvillian $\mathcal{L}=[H,\cdot]$ on $\partial_{\lambda}H$ Claeys et al. (2019):
@@
Notably, only odd orders of nested commutators enter into this ansatz. Here $\ell$ controls the order (and therefore the locality) of the expansion, and $\alpha_{k}$ are the variational parameters found by minimizing the action (4). Formally this ansatz is exact in the limit $\ell\to\infty$ but in practice we want to restrict it to small values of $\ell$. This process can by made more numerically stable by using the Lanczos algorithm in the operator basis leading to an AGP expansion in an orthonormal set of Krylov operators $O_{k}$ (see Refs. Takahashi and del Campo, 2023; Bhattacharjee, 2023; Sels and Polkovnikov, 2023 for details). We give a detailed description of our Krylov space construction of the AGP in Appendix A.
@@ III Improving CD driving with extra controls
The main goal of this section and of the whole paper is to find a systematic approach to improving performance of approximate CD driving via adding extra controls to the Hamiltonian. We test these by applying them to prepare nontrivial GHZ entangled states. We will begin by focusing on the infinite speed limit $|\dot{\lambda}|\to\infty$, where although the $H$ term in (1) is included in the simulations, the evolution according to $H_{\rm CD}$ is dominated by the counterdiabatic term $\dot{\lambda}A_{\lambda}$. In III.3, we discuss the improvement from the extra controls in the regime where $\dot{\lambda}$ is finite, when both $H$ and $A_{\lambda}$ play a meaningful role in the dynamics.
@@
with periodic boundary conditions. We will start from the situation where there is a small but finite longitudinal field $0<h_{z}\ll 1$ and finite $h_{x}<1$. The first condition breaks the $Z_{2}$ symmetry of the model such that the ferromagnetic ground state for $|h_{x}|<1$ is polarized in the positive $z$-direction and the second condition implies that during the annealing protocol the system crosses a quantum phase transition at $h_{x}=1$ Sachdev (2011). At this point the gap closes and the AGP becomes an operator with an infinite range del Campo, Rams, and Zurek (2012); Damski (2014); Kolodrubetz et al. (2017) such that local CD protocols become very inefficient.
Figure 2: A comparison between the results when performing local CD driving on the “naive” original path in red, which passes close by the critical point at $(h_{x},h_{z})=(1,0)$, and the augmented path in blue, where the critical point is avoided. The final Hamiltonian $H_{0}$ is given by Eq. (6) with $h_{x}=0.7,\;h_{z}=0.01$. The circles and squares show the results using $\ell=1$ and $\ell=2$ in Eq. (5) respectively. We restrict to $\beta\leq 3$ as defined in Eq. (7) to ensure that we are not finding optimal paths by taking $\beta$ large and effectively rescaling time.
@@
$\tilde{H}(\lambda)=H(\lambda)+\beta\sin(\pi\lambda)\sum_{i}\sigma_{i}^{z},$ (7)
@@ III.1 Formulation of General Problem
where $H(\lambda$) is given by Eq. (6) and the $\beta^{(n)}(\lambda)$ are some smooth functions satisfying boundary conditions $\beta^{(n)}(\lambda=0)=\beta^{(n)}(\lambda=1)=0$ so that the initial and final eigenstates of the annealing Hamiltonian are unchanged. We refer to $H_{c}$ as an extra control Hamiltonian. In the previous example, it was just an additional global $z$ field, i.e. $H_{c}=\sum_{i}\sigma_{i}^{z}$. We stress that in the infinite speed limit the dynamics is determined exclusively by the AGP, so the annealing Hamiltonian plays a purely auxiliary role like e.g. in Ref. Ljubotina et al., 2022.
We then perform local CD driving for the augmented Hamiltonian $\tilde{H}(\lambda)$, hence evolving according to $\tilde{H}_{CD}(\lambda)$.The goal is to choose both $\beta^{(n)}(\lambda)$ and $H_{c}^{(n)}$ to maximize the fidelity of the final state. The $\beta^{(n)}(\lambda)$ are fixed for a given evolution, and the variational parameters $\alpha_{k}$ of the AGP are determined locally at each time step. Finding the optimal control functions $\beta^{(n)}(\lambda)$ is the subject of quantum optimal control Glaser et al. (2015) or similar methods Bukov et al. (2018), which is not the main focus of this work. We pick the $\beta^{(n)}(\lambda)$ by a very simple optimization, following earlier work Čepaitė et al. (2023), writing it as a single harmonic term
@@
and choosing $\beta^{(n)}$ to maximize the fidelity of the final state. This is done by maximizing the fidelity
@@
where |$\ket{\psi_{evolved}(\beta)}$ is obtained by evolving the initial state by the augmented CD Hamiltonian $\tilde{H}_{CD}$. We note that the fidelities can be even further improved by adding additional couplings $\beta_{k}^{(n)}$ corresponding to $k$-th harmonic of $\pi\lambda$epaitė et al. (2023). We limit ourselves only to the first one as we want to focus on the question of finding optimal $H_{c}^{(n)}$.
@@ III.2 Ansatz for Extra Controls
To proceed, let us consider a state preparation problem similar to the previous one, but with $h_{x}=h_{z}=0$, i.e. without breaking the $Z_{2}$-symmetry. The ferromagnetic ground state of this model is now a GHZ state Greenberger, Horne, and Zeilinger (2007). It is defined as an equal superposition of all spins up and all spins down. This adiabatic state preparation is encoded in the Hamiltonian
@@
with the corresponding weights $\beta^{(n)}(\lambda)$ as in Eq. (8). By adding such terms we clearly expand the variational manifold allowing one to gradually increase the locality of both $\tilde{H}(\lambda)$ and $A_{\lambda}$. Moreover such terms can be generated using Floquet pulse sequences containing only $H_{0}$ and $H_{1}$ and is routinely done in the NMR literature Farrar and Becker (1971). A detailed description of how to implement a version of our protocol with Floquet pulses is given in Appendix B. A very similar approach is described in Claeys et al., 2019. As we show below, for the examples we analyze here this idea leads to the dramatic improvement of local CD protocols.
It is easy to check that for the Ising model the new terms (i.e. terms which are not originally present in $H_{0}$) which appear by computing the commutators in Eq. (12) are $YY$ and $ZXZ$, where we use a common short-hand notation: $YY=\sum_{j}\sigma_{j}^{y}\sigma_{j+1}^{y}$, $ZXZ=\sum_{j}\sigma_{j}^{z}\sigma_{j+1}^{x}\sigma_{j+2}^{z}$. We can thus select $H_{c}^{(1)}=YY$ and $H_{c}^{(2)}=ZXZ$. In this way the augmented Hamiltonian (8) will cause the state to follow a genuinely new path. The resulting improvement of the final state fidelity is plotted in Figure 3. For a more detailed discussion of the optimization procedure see Appendix C.
Figure 3: Improvement in the final state fidelity when prepared with local counterdiabatic driving. In red, we perform local CD driving along the original path given by Eq. (6), whereas in blue it is performed along a path where we augment Eq. (6) by $H_{c}^{(1)}=YY$ and $H_{c}^{(2)}=ZXZ$ as in (12). The circles and squares of red and blue lines correspond to $\ell=1$ and $\ell=2$ in Eq. (5) respectively.
@@
III.3 Finite-Time Protocol Performance
Up to this point, we have only considered the case where $|\dot{\lambda}|\to\infty$ and the dynamics are controlled entirely by the AGP $A_{\lambda}$ in Eq. (1). There is also the trivial adiabatic limit, where $|\dot{\lambda}|\to 0$ and the eigenstates are transported perfectly with the AGP playing no role. But in between these, there is a regime where both the Hamiltonian and the AGP play an important role in the counterdiabatic dynamics.
For the same annealing problem defined in Eq. (11), we plot the final state fidelity when we perform CD driving with different control schemes in Figure 4. From this, one can see that different extra control schemes offer improvement in the finite time regime, in addition to the $|\dot{\lambda}|\to\infty$ ($\tau\to 0$) limit discussed earlier. In particular, using the recipe for extra controls given in Eq. (12), one can prepare states with unit fidelity in a time reduced by nearly three orders of magnitude.
Figure 4: The final state fidelity for the annealing problem of Eq. (11) with $N=8$, for different total protocol times. The final state fidelity is improved not only in the “fast” limit with $\tau\to 0$, but for any intermediate protocol duration up to the adiabatic limit. Particularly noteworthy is that even local controls are sufficient to reduce the threshold for adiabatic evolution by several orders of magnitude.
IV Preparing GHZ state in a Long range model
IV.1 Fidelity Gain from Augmented Hamiltonian
We now move to preparation of the GHZ state in longer range Ising Hamiltonians, which are relevant to cold atom/trapped ion systems Koffel, Lewenstein, and Tagliacozzo (2012); Jaschke et al. (2017); Adelhardt et al. (2020). We define the annealing Hamiltonian for a long-range Ising model: